OpenStudy (anonymous):

Im clueless on these questions!!! PLEASE HELP!!! (1) Solve by Substitution: 2x + y = 0 x – 3y = -7 (2) Solve by Elimination: 3x + 4y = 1 5x – 3y = 21

OpenStudy (anonymous):

I will fan and give medal!!!

OpenStudy (anonymous):

Alright. I will io step by step. Ok?

OpenStudy (anonymous):

go*

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

make one a y=mx+b equation and put in place of the other equations y

OpenStudy (anonymous):

(1) Solve by Substitution: 2x + y = 0 x – 3y = -7 you want one equation to equal either x or y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so out of: 2x + y = 0 x – 3y = -7 it would be easier to get x by itself in the second question x – 3y = -7 add 3y to both sides x= 3y -7

OpenStudy (ciel21):

simple make one a y= m x + b equation and put into the place of the other equations y

OpenStudy (anonymous):

ok @WinterRaine ... i'm following.

OpenStudy (anonymous):

so you substitute x= 3y -7 into the other equation 2x + y = 0 2(3y -7) +y =0

OpenStudy (anonymous):

in the 2nd equation, wouldn't it be 2(3y -7) + x =0 ?? instead of y?

OpenStudy (anonymous):

your two equations look like this: x – 3y = -7 (x= 3y -7) & 2x + y = 0

OpenStudy (anonymous):

Yes, I understand that.

OpenStudy (anonymous):

so when you substitute x – 3y = -7 into the other equation, you just put it where the x is so 2(3y -7) +y =0

OpenStudy (anonymous):

ohhhhhh ok! please continue (:

OpenStudy (anonymous):

alright so you solve the equation for y 2(3y -7) +y =0 distribute 6y - 14 + y=0 combine like terms 7y -14 =0 get the y by itself (add 14 to both sides) 7y = 14 divide by 7 on both sides y =2

OpenStudy (anonymous):

ok. now we have to solve for x now right?

OpenStudy (anonymous):

yeah. so substitute y=2 into x= 3y -7

OpenStudy (anonymous):

x = -1

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

((((((((:

OpenStudy (anonymous):

Thank you soooo much. Mind doing the other one with me?

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

(2) Solve by Elimination: 3x + 4y = 1 5x – 3y = 21

OpenStudy (anonymous):

we'll use this equation to start off: 3x + 4y = 1 we'll get x by itself for this one subtract 4y from both sides 3x = -4y +1

OpenStudy (anonymous):

3x = -4y +1 divide by 3 on both sides x = -4/3x + 1/3

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

so put x = -4/3y + 1/3 into 5x – 3y = 21 5(-4/3y + 1/3) -3y =21

OpenStudy (anonymous):

-6 2/3y + 5/3 - 3y= 21

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

OR -20/3y + 5/3 -3y= 21 combine like terms -29/3y + 5/3 = 21 get y by itself (subtract 5/3) -29/3y = 19.34

OpenStudy (anonymous):

I'm sorry-- I have to go. @Easyaspi314 should be able to help you along and explain it better than I have (:

OpenStudy (anonymous):

Good luck & I hope you understand substitution better!

OpenStudy (anonymous):

ok thanks. @WinterRaine ... @RosieS12 could you finish where he left off?

OpenStudy (anonymous):

so which part did u not understand?